Graham, Craig, Neil
Minor correction - Alastair, rather than George, commented on 1 month's notice for power shutdown.
Report from Computing Executive Group
Reports from units.
The first stage has now been deployed. The conduits are being run manually for now but will be automated shortly.
Stephen explained that the fix for the memory leak problem associated with the recent perl package update is not trivial. It does not appear to be a problem with f13.
A deadline of the end of August for all SL5.5 server reboots was agreed. It was also noted that f13 will not be available for the 260, 270 and hp530. We do still have a few of these machines in service. They will need to either be brought out of service completely, replaced or become self-managed machines.
Ian noted that he had spotted fibre switches resetting themselves when these reboots took place.
Topics for discussion
This was not discussed at any length (see Services report) but it was noted that we have possibly not seen the issue for a few weeks which might co-incide with the update to openafs 1.4.12.
Alison to confirm with Sheila on server delivery status. There have been significant delays and George Reid has been made aware.
The next meeting should be on Wednesday 25th August in room 4.31.
Please contact us with any comments or corrections.
Unless explicitly stated otherwise, all material is copyright The University of Edinburgh